One train leaves a station heading due west. Two hours later a second train leaves the same station heading due east. The second train is traveling 15 mi/h faster than the first. Six hours after the second train leaves, the two trains are 580 miles apart. Find the rate at which each train is traveling.
Answer by mananth(16949)

Train A speed = x mph
Train B speed = x + 15 mph
Train A has left 2 hours early
Train A has traveled 2 x miles before Train B starts
The trains are travelling in opposite directions
so their combined speed = 2 x+ 15 mph
Train B has traveled 6 hours
Diatance 580
2x+6( 2 x + 15 ) = 580
2x + 12x + 90 = 580
14 x = 580 -90
14 x = 490
x = 35 mph Train A speed
50 mph Train B speed